### v3

@@ Abstract
Camouflaged object detection and segmentation is a new and challenging research topic in computer vision. There is a serious issue of lacking data on concealed objects such as camouflaged animals in natural scenes. In this paper, we address the problem of few-shot learning for camouflaged object detection and segmentation. To this end, we first collect a new dataset, CAMO-FS, for the benchmark. We then propose a novel method to efficiently detect and segment the camouflaged objects in the images. In particular, we introduce the instance triplet loss and the instance memory storage. The extensive experiments demonstrated that our proposed method achieves state-of-the-art performance on the newly collected dataset.
@@ I Introduction
Camouflage is a defense mechanism that animals use to conceal their appearance by blending in with their environment [1]. Autonomously detecting camouflaged animals is helpful in various applications, e.g., search-and-rescue missions [2]; wild species discovery and preservation activities [2]; and media forensics (manipulated image/video detection and segmentation [3]). Although image segmentation methods have been proposed for a long time, general detectors cannot deal with camouflaged animals [4, 5, 6, 7]. The detectors initially developed for camouflage detection [8, 9, 10, 11, 12, 13, 14, 15], which use handcrafted low-level features, are effective only for images with a simple and uniform background. More recently developed deep learning-based detectors [2, 16] for camouflaged object segmentation rely on large-scale data.
@@
Dataset Year Venue Type #Annot. Camo. Img. #Meta- Cat. #Obj. Cat. #Ins. or #Obj. per Img. Bbox. GT Obj. Mask GT Ins. Mask GT Few-shot CamouflagedAnimals [17] 2016 ECCV Video 181 - 6 1.238 $\times$ $\times$ MoCA [18] 2020 ACCV Video 7,617 - 67 1.000 $\times$ $\times$ $\times$ CHAMELEON [19] 2018 - Image 76 - - 1.000 $\times$ $\times$ $\times$ CAMO [2] 2019 CVIU Image 1,250 2 8 1.000 $\times$ $\times$ $\times$ COD [16] 2020 CVPR Image 5,066 5 69 1.171 $\times$ NC4K [20] 2021 CVPR Image 4,121 5 69 1.171 $\times$ CAMO++ [21] 2022 TIP Image 2,695 10 47 1.171 $\times$ CAMO-FS (Ours) 2023 - Image 2,858 10 47 1.172
With the approach of few-shot learning, we can perform machine learning tasks with given limited data. A few-shot method requires two stages of processing: (1) one base phase training for the model to gain concept knowledge of general domains with abundant data, and then (2) performing a novel phase that can do the specific task on few-shot data. In the case of the camouflaged object detection and segmentation task, we leverage few-shot learning in the concept of camouflaged animals which is rare and hard to find in the wild. Thus, with limited data on camouflaged objects, the models can still well handle the given tasks. However, none of the aforementioned publications targets few-shot camouflaged object detection and segmentation despite its practical applications. The task of segmentation supports better to identify camouflaged objects in terms of specifying which pixels in the images conceal the objects in comparison to classification and detection. The research on camouflaged animals suffers due to the lack of data. There are not many object classes with rare instances captured in photos. Therefore, in this paper, we would like to address few-shot learning with camouflaged animals.
Our contributions in this work are two-fold:
@@
  • Second, we propose a framework to efficiently detect and segment camouflaged instances, named after FS-CDIS, given a small shot of training data for novel classes utilizing the idea of instance triplet loss and instance memory storage.
@@ II-A Camouflage Research
Given any region (i.e. bounding boxes or polygon masks) presented for an object of interest (i.e. animals or artificial objects) in an image and then they tend to be classified as background, contents in that region can be qualified as camouflaged objects. Thus, a camouflaged object is defined as a set of bounding boxes or camouflaged pixels in an image without any further detailed information such as the number of objects or the semantic meaning [2]. Although tasks related to camouflaged animals are performed in a wide range of applications, this research field has not been well explored in the literature, especially few-shot learning which is practically suitable to the context of scare data as camouflaged animals.
Binary camouflage segmentation. Prior to the advancement of deep neural networks, most of the work exploits identical regions between camouflaged regions and the background by handcrafted or low-level features, specifically based on external characteristics (e.g., color, shape, orientation, and brightness). Particularly, early camouflage detection works had attention on the foreground region even when some of its texture was similar to the background [8, 22, 9, 10]. The foreground was distinguishable from the background via simple features, such as color, intensity, shape, orientation, and edge [23, 24, 25, 22, 26]. A few methods  [11, 27, 12, 13, 14, 15] based on handcrafted low-level features have been proposed for tackling the problem of camouflage detection. However, they are effective only for images with a simple and uniform background. Thus, their performances are unsatisfactory in camouflaged object segmentation due to the substantial similarity between the foreground and the background.
@@
Camouflage datasets. CamouflagedAnimals [17] and CHAMELEON [19] were the first two camouflage datasets with mask annotations. The two datasets do not contain enough images to train deep learning methods. Le et al. [2] created the CAMO dataset, the first camouflage dataset with more than $1,000$ annotated images. It contains $1,250$ annotated images, which is a limited number of samples to train and evaluate deep learning methods. Then, Fan et al. [16] collected the COD dataset, which comprises $10,000$ images (both camouflage and non-camouflage) divided into 5 meta-categories. However, they annotated only $5,066$ camouflage images. Lamdouar et al. [18] recently developed the MoCA dataset for the camouflage object detection task; it contains only bounding box ground truths. Hence, these datasets limit their annotations at binary ground truth datasets which have a shortage of intensive annotations for multi-task camouflage problems. CAMO++ [21] is different from the above-mentioned dataset, CAMO++ provides a benchmark for camouflaged instance segmentation with more comprehensive annotations and diverse meta-categories of 10. The dataset comprises $5,500$ images with superiority over other datasets on instances including $32,756$ instances for both camo and non-camo objects.
@@ II-B Few-shot Learning
Few-shot object segmentation (FSOS). Recently, the field of few-shot segmentation gained attention from the community. As mentioned above, the first work Meta RCNN originated from Mask RCNN, therefore, Meta RCNN simultaneously performed detection and segmentation. Liu et al. [56] utilized a cross-reference network for generic image segmentation. The authors proposed a cross-reference mechanism and a mask refinement module to specifically support the task of segmentation. Before, Dong et al. [57] proposed a prototype learning component in a framework of semantic segmentation that learned to take discriminative information from features to help segment objects better. Also, Wang et al. [58] introduced a prototype align method that learns class-specific prototype representations from a few image samples to perform segmentation over the query images. Lately, Liu et al. proposed a dynamic prototype convolution network to address few-shot semantic segmentation. The work of [59] proposed context-aware prototype learning. [60] introduced generative models approach for this task. Recently, Nguyen et al. [61] came up with iFS-RCNN, an instance segmenter via an incremental approach. Gao et al. [62] proposed the DCFS framework, an effective decoupling classifier that boosted the performance of object detection and segmentation heads. Han et al. [63] suggested a reference twice transformer-based framework (RefT) to enhance features in segmentation tasks. Also in the transformer approach, Wang et al. [64] introduced DTN to directly segment the target object instances from arbitrary categories given reference images. In common, these aforementioned methods focus on generic objects.
III Proposed Method
@@
III-A CAMO-FS Benchmark Dataset
#Instances Ratio (%) #Images 1 90.5 2581 2 1.05 190 3 1.79 51 3$+$ 6.66 30
@@ III-A CAMO-FS Benchmark Dataset
CAMO-FS Dataset. We leverage the available CAMO++ to build our CAMO-FS dataset. In this way, we inherit the biology taxonomic and vision taxonomic structure of CAMO++ which helps us to reduce the burden of data collection. Table I provides an overview of previous works done on camouflage, which is mentioned in the related work, and our proposed CAMO-FS in terms of main characteristics. We exploit the diversity of CAMO++ by its 10 meta-categories to build up the few-shot concept for instance segmentation. To this end, our CAMO-FS not only keeps a good ratio of instances per image of 1.172 but also contributes as the very first dataset specific for few-shot research on camouflaged animals. Note that the large amount of images in some datasets does not mean they are all camouflaged images.
However, imbalanced data and a shortage of the number of images of some classes inherently exist in CAMO++ posing problems of evaluation for few-shot tasks. Particularly, there are 11 classes (e.g. Camel, Dolphin, Elephant, Horse, Kangaroo, Monkey, Penguin, Bat, Bear, Squirrel, and Rhino) having a shortage of images that are needed to train a few-shot model. Hence, we hardly perform training or testing on these classes. As a result, we collect more data for these classes with 163 total images corresponding to 181 instances (an average of 15-16 instances per class). We also remove images with mistakes in the original dataset. The statistics of collected data are shown in Table III. By gathering more camouflaged animals and combining them with the CAMO++ dataset, we conduct our CAMO-FS dataset for few-shot camouflaged animal detection and segmentation with 2,858 total images corresponding to 3,342 instances. Figure 1 shows exemplary images with mask annotations from our proposed CAMO-FS.
To effectively create the data for the few-shot problem, we get $M$ instances from the CAMO-FS dataset to create training sets (in our setup, $M=5$) and use the remaining instances for testing. We only remove some objects of the higher-level training set if they exist to create the other few-shot settings. For example, we get all elements to generate 5-shot training data and discard 2 in 5 objects to make a 3-shot one. In this way, the 5-shot benchmark contains objects of the 3-shot dataset and the 3-shot setting contains the objects of the 2-shot one.
@@ III-B General Framework
Figure 5: Illustration of the instance triplet loss on each feature proposal.
@@
Figure 6: Visualization of memory loss for region proposal.
@@ III-C2 Instance memory storage
TABLE IV: State-of-the-art comparison on CAMO-FS dataset among the baseline model of MTFA [70], Mask ${}^{\dagger}$ [40], iFS-RCNN [61], and our proposed methods FS-CDIS with instance triplet loss (-ITL) and instance memory storage (-IMS). Our performance improves over the utilized baselines.
Model Novel AP Method Baseline Instance Segmentation Object Detection 1 2 3 5 Avg. 1 2 3 5 Avg. MTFA [70] COCO-80 ResNet-50 2.48 6.67 5.81 6.40 5.34 1.98 6.47 5.82 6.17 5.11 ${}^{\dagger}$ [40] 4.08 6.79 6.90 8.29 6.52 2.82 5.09 5.46 6.18 4.89 iFS-RCNN [61] 4.17 6.26 5.73 6.38 5.64 3.92 6.06 5.47 6.60 5.51 MTFA [70] COCO-80 ResNet-101 3.66 6.21 6.16 5.95 5.50 2.93 5.90 5.84 5.84 5.13 ${}^{\dagger}$ [40] 4.39 7.69 7.94 10.09 7.53 3.03 5.80 6.20 7.79 5.71 iFS-RCNN [61] 4.27 6.55 6.07 7.80 6.17 3.79 6.28 6.01 8.08 6.04 Our performance (normal-$\uparrow$) FS-CDIS-ITL ResNet-101 4.46 5.57 6.41 8.48 6.23 4.04 7.28 7.49 9.76 7.14 FS-CDIS-IMS MTFA 5.46 6.95 7.36 9.61 7.35 4.50 6.95 7.55 10.36 7.34 FS-CDIS-ITL ResNet-101 5.73 7.97 8.52 9.92 8.04 5.08 7.56 7.85 9.67 7.34 FS-CDIS-IMS M-RCNN 5.52 7.84 8.65 9.82 7.96 4.92 7.39 7.96 9.52 7.45 FS-CDIS-ITL ResNet-101 5.35 6.01 7.80 9.35 7.13 4.71 5.66 7.10 10.36 6.96 FS-CDIS-IMS iFS-RCNN 2.99 6.83 6.14 9.03 6.25 2.74 6.39 5.94 8.44 5.88 ${}^{\dagger}$ is Mask R-CNN [40] with sigmoid classifier.

### v4

@@ Abstract
Camouflaged object detection and segmentation is a new and challenging research topic in computer vision. There is a serious issue of lacking data on concealed objects such as camouflaged animals in natural scenes. In this paper, we address the problem of few-shot learning for camouflaged object detection and segmentation. To this end, we first collect a new dataset, CAMO-FS, for the benchmark. As camouflaged instances are challenging to recognize due to their similarity compared to the surroundings, we guide our models to obtain camouflaged features that highly distinguish the instances from the background. In this work, we propose FS-CDIS, a framework to efficiently detect and segment camouflaged instances via two loss functions contributing to the training process. Firstly, the instance triplet loss with the characteristic of differentiating the anchor, which is the mean of all camouflaged foreground points, and the background points are employed to work at the instance level. Secondly, to consolidate the generalization at the class level, we present instance memory storage with the scope of storing camouflaged features of the same category, allowing the model to capture further class-level information during the learning process. The extensive experiments demonstrated that our proposed method achieves state-of-the-art performance on the newly collected dataset. Code is available at https://github.com/danhntd/FS-CDIS.
@@ I Introduction
Camouflage is a defense mechanism that animals use to conceal their appearance by blending in with their environment [1]. Autonomously detecting camouflaged animals is helpful in various applications, e.g., search-and-rescue missions [2]; wild species discovery and preservation activities [2]; and media forensics [3, 4, 5, 6](manipulated image/video detection and segmentation [7]). By leveraging camouflage recognition at object detection or instance segmentation level autonomously, these practical applications can be done with minor efforts from humans while maintaining work performance. To be specific, utilizing a drone flying around the mountainous area to collect images and videos for a system to detect and segment the target object in danger is more effective than sending a group of humans manually scanning the zone. By this means, this process can support biological scientists in identifying and preserving endangered species effectively. Further related applications can be considered in different important fields including healthcare, agriculture, or military, where exist the concept of finding objects with camouflaged features. Indeed, camouflage detection and segmentation tasks can provide further applications such as assisting doctors in medical imaging understanding, supporting modern farmers in managing the vast fields of crops via visual information or even detecting hidden enemies in nature. These applications can be potential via the development of camouflaged research via image understanding at detailed levels of object detection and instance segmentation.
@@
Dataset Year Venue Type #Annot. Camo. Img. #Meta- Cat. #Obj. Cat. #Ins. or #Obj. per Img. Bbox. GT Obj. Mask GT Ins. Mask GT Few-shot CamouflagedAnimals [8] 2016 ECCV Video 181 - 6 1.238 $\times$ $\times$ MoCA [9] 2020 ACCV Video 7,617 - 67 1.000 $\times$ $\times$ $\times$ CHAMELEON [10] 2018 - Image 76 - - 1.000 $\times$ $\times$ $\times$ CAMO [2] 2019 CVIU Image 1,250 2 8 1.000 $\times$ $\times$ $\times$ COD [11] 2020 CVPR Image 5,066 5 69 1.171 $\times$ NC4K [12] 2021 CVPR Image 4,121 5 69 1.171 $\times$ CAMO++ [13] 2022 TIP Image 2,695 10 47 1.171 $\times$ CAMO-FS (Ours) 2024 IEEE ACCESS Image 2,852 10 47 1.172
Although image segmentation methods have been proposed for a long time, general detectors cannot deal with camouflaged animals [14, 15, 16, 17] due to their specific camouflaged features. The detectors initially developed for camouflage detection [18, 19, 20, 21, 22, 23, 24, 25], which use handcrafted low-level features, are effective only for images with a simple and uniform background. More recently developed deep learning-based detectors [2, 26, 27, 11, 28, 29, 30, 12, 10] for camouflaged object segmentation. Most previous methods are trained on large-scale datasets to perform computer vision tasks. Nevertheless, building such standard datasets for camouflaged objects is labor-intensive due to the ambiguity between the objects and their backgrounds, which leads to more time and cost in the labeling procedure. To address the problem, we consider the camouflaged object detection and instance segmentation under the few-shot learning which shows potential results when utilizing limited labeled samples to classify [31, 32, 33, 34, 35, 36, 37, 38], detect [39, 40, 41, 42, 43, 44, 45, 46, 47, 48], or even segment [49, 50, 51, 52, 53] new objects. However, to the best of our knowledge, there is a lack of camouflaged datasets supporting few-shot learning in camouflaged research. Therefore, we introduce a new benchmark, CAMO-FS, for camouflaged few-shot object detection and instance segmentation under the few-shot settings. The new benchmark is mainly reconstructed from the CAMO++ dataset [13] due to its diversity. The new benchmark consists of $197$ camouflaged images for training and $2,655$ camouflaged images for performance evaluation as described in Table I.
There are several approaches for few-shot learning. Beginning with few-shot classification (FSL), many works are based on meta-learning [31, 32, 33, 34, 54] or transfer-learning [35, 36, 37, 38] approaches to leverage a few labeled data to classify new objects and achieve incredible results. The former approaches compute the similarity between query and support images to pinpoint novel objects while the latter involves utilizing knowledge from the source domain to adapt a different but related target domain. Fueled by these successes, most existing works on few-shot object detection (FSOD) and few-shot segmentation (FSS) which are recently developed to tackle the problem through meta-learning [42, 43, 44, 45, 46, 47, 48, 55, 56, 57, 51, 52, 53] and transfer-learning [49, 39, 40, 41, 58, 50, 59] methods. Nonetheless, such methods focus on the general domain and thus fail to generate effective features for camouflaged objects due to the ambiguity between backgrounds and foregrounds. To overcome the specific issue of camouflage objects, we propose a novel framework for few-shot camouflaged object detection and instance segmentation, dubbed FS-CDIS, which is based on the transfer-learning approach. The model is trained on two stages of processing: (1) one base phase training for the model to gain concept knowledge of general domains with abundant data, and then (2) performing a novel phase that can do the specific task on the few-shot data. To be specific, in the base training stage, we train our model on generic object detection and instance segmentation dataset (e.g. COCO [60]) and focus on improving the model in the novel fine-tuning stage. Because of the similarity between camouflaged objects and their surroundings, we aim to guide our few-shot models to obtain camouflaged features that highly distinguish the instances from the background. To achieve that target, we introduce two loss functions contributing directly to the novel fine-tuning process. Firstly, the instance triplet loss with the characteristic of differentiating the anchor, which is the mean of all camouflaged foreground points, and the background points are employed to work at the instance level. Secondly, to consolidate the generalization at the class level, we present instance memory storage with the scope of storing camouflaged features of the same category, allowing the model to capture further class-level information during the learning process.
To summarize, our contributions in this work are two-fold:
@@
  • Second, we propose a framework to detect and segment camouflaged instances efficiently, named after FS-CDIS, given a small shot of training data for novel classes utilizing the idea of instance triplet loss and instance memory storage.
@@ II-A Camouflage Research
Given any region (i.e. bounding boxes or polygon masks) presented for an object of interest (i.e. animals or artificial objects) in an image and then they tend to be classified as background, contents in that region can be qualified as camouflaged objects. Thus, a camouflaged object is defined as a set of bounding boxes or camouflaged pixels in an image without any further detailed information such as the number of objects or the semantic meaning [2]. Although tasks related to camouflaged animals can be performed in a wide range of applications such as security systems [3, 4], pollution detection [5], watermark detection [6], this research field has not been well explored in the literature, especially few-shot learning which is practically suitable to the context of scare data as camouflaged animals.
Binary camouflage segmentation. Prior to the advancement of deep neural networks, most of the work exploits identical regions between camouflaged regions and the background by handcrafted or low-level features, specifically based on external characteristics (e.g., color, shape, orientation, and brightness). Particularly, early camouflage detection works had attention on the foreground region even when some of its texture was similar to the background [18, 19, 19, 20]. The foreground was distinguishable from the background via simple features, such as color, intensity, shape, orientation, and edge [61, 18, 62, 19, 20]. A few methods  [21, 22, 23, 24, 25] based on handcrafted low-level features have been proposed for tackling the problem of camouflage detection. However, they are effective only for images with a simple and uniform background. Thus, their performances are unsatisfactory in camouflaged object segmentation due to the substantial similarity between the foreground and the background.
@@
Camouflage datasets. CamouflagedAnimals [8] and CHAMELEON [10] were the first two camouflage datasets with mask annotations. The two datasets do not contain enough images to train deep learning methods. Le et al. [2] created the CAMO dataset, the first camouflage dataset with more than $1,000$ annotated images. It contains $1,250$ annotated images, which is a limited number of samples to train and evaluate deep learning methods. Then, Fan et al. [11] collected the COD dataset, which comprises $10,000$ images (both camouflage and non-camouflage) divided into 5 meta-categories. However, they annotated only $5,066$ camouflage images. Lamdouar et al. [9] recently developed the MoCA dataset for the camouflage object detection task; it contains only bounding box ground truths. Hence, these datasets limit their annotations at binary ground truth datasets which have a shortage of intensive annotations for multi-task camouflage problems. CAMO++ [13] is different from the aforementioned datasets providing a benchmark for camouflaged instance segmentation with more comprehensive annotations and diverse meta-categories of 10. The dataset comprises $5,500$ images with superiority over other datasets on instances including $32,756$ instances for both camo and non-camo objects. Different from the existing work, we address camouflage research under the few-shot learning concept to detect objects and segment camouflaged instances. Therefore, we introduce a new benchmark, dubbed CAMO-FS, to support the evaluation process of this specific task. Accordingly, our CAMO-FS comprises $2,852$ images as a result of the inheritance from CAMO++  [13] and our further collection. This specific dataset serves $93,1\%$ of the annotated images for the evaluation process while the rest few samples are provided for training (i.e. $197$ images for training and $2,655$ images for testing).
@@ II-B Few-shot Learning
Few-shot object segmentation (FSS). Recently, the field of few-shot segmentation gained attention from the community. As mentioned above, the first work Meta RCNN originated from Mask RCNN, therefore, Meta RCNN simultaneously performed detection and segmentation. Liu et al. [74] utilized a cross-reference network for generic image segmentation. The authors proposed a cross-reference mechanism and a mask refinement module to specifically support the task of segmentation. Before, Dong et al. [75] proposed a prototype learning component in a framework of semantic segmentation that learned to take discriminative information from features to help segment objects better. Also, Wang et al. [76] introduced a prototype align method that learns class-specific prototype representations from a few image samples to perform segmentation over the query images. Lately, Liu et al. proposed a dynamic prototype convolution network to address few-shot semantic segmentation. The work of [77] proposed context-aware prototype learning. [78] introduced generative models approach for this task. Recently, Nguyen et al. [79] came up with iFS-RCNN, an instance segmenter via an incremental approach. Gao et al. [80] proposed the DCFS framework, an effective decoupling classifier that boosted the performance of object detection and segmentation heads. Han et al. [81] suggested a reference twice transformer-based framework (RefT) to enhance features in segmentation tasks. Also in the transformer approach, Wang et al. [82] introduced DTN to directly segment the target object instances from arbitrary categories given reference images.
In common, these aforementioned methods of the two approaches including FSOD and FSS mostly focus on generic objects, which cannot create effective distinguished features and fail to recognize camouflaged objects. In our case, our proposed methods aim at highlighting the differences between backgrounds and foregrounds which we considered as the key feature to detect or segment camouflaged objects. Furthermore, our proposed approaches contribute directly to the training process of such models via loss functions.
#Instances Ratio (%) #Images 1 85.55 2440 2 8.31 237 3 3.44 98 3$+$ 2.70 77
@@
Figure 2: Hierarchical taxonomic structure of our CAMO-FS dataset.
III Proposed Method
III-A CAMO-FS Benchmark Dataset
@@
Figure 3: The distribution of CAMO-FS dataset. The categories are sorted.
CAMO-FS Dataset. In general, there are three steps to construct the CAMO-FS dataset: inheritance, collection, data splitting. In the inheritance step, we leverage the available CAMO++ to build our CAMO-FS dataset. In this way, we inherit the biology taxonomic and vision taxonomic structure of CAMO++ which helps us to reduce the burden of data collection. Table I provides an overview of previous works done on camouflage, which is mentioned in the related work, and our proposed CAMO-FS in terms of main characteristics. We exploit the diversity of CAMO++ by its 10 meta-categories to build up the few-shot concept for instance segmentation. To this end, our CAMO-FS not only keeps a good ratio of instances per image of 1.172 but also contributes as the very first dataset specific for few-shot research on camouflaged animals. Note that the large amount of images in some datasets does not mean they are all camouflaged images.
In the collection step, as CAMO++ faces issues such as imbalanced data and a shortage of the number of images of some classes, which cause evaluation problems for few-shot tasks. Particularly, there are 11 classes (e.g. Camel, Dolphin, Elephant, Horse, Kangaroo, Monkey, Penguin, Bat, Bear, Squirrel, and Rhino) having a shortage of images that are needed to train a few-shot model. Hence, we hardly perform training or testing on these classes. As a result, we manually collect more data for these classes with 163 total images corresponding to 181 instances (an average of 15-16 instances per class) on Google Image Search Engine with their class names as the search query. We also remove images with mistakes in the original dataset. The statistics of collected data are shown in Table III. By gathering more camouflaged animals and combining them with the CAMO++ dataset, we conduct our CAMO-FS dataset for few-shot camouflaged animal detection and segmentation with $2,852$ total images corresponding to $3,342$ instances. Figure 2 shows the vision taxonomic structure of coarse-grained and corresponding fine-grained classes and illustrates the ratios of 10 coarse-grained classes in our proposed CAMO-FS dataset. We also show the distribution of 47 camouflaged classes in Figure 3, which indicates that CAMO-FS is a diverse and long-tailed dataset. Figure 1 shows exemplary images with mask annotations from our proposed CAMO-FS.
@@
In the data splitting step, to effectively create the data for the few-shot problem, we randomly get $M$ instances for each camouflaged class from the CAMO-FS dataset to create training sets. In our setup, $M$ equals $5$ for $47$ classes and thus there are $197$ training images containing $235$ camouflaged instances and the remaining $2,655$ images with $3,107$ instances are used for testing. We only remove some objects of the higher-level training set if they exist to create the other few-shot settings. For example, we get all elements to generate 5-shot training data and discard 2 in 5 objects to make a 3-shot one. In this way, the 5-shot benchmark contains objects of the 3-shot dataset and the 3-shot setting contains the objects of the 2-shot one.
@@ III-B General Framework
Figure 7: Visualization of instance triplet loss and instance memory loss for region proposal.
@@ III-C2 Instance memory storage
TABLE IV: State-of-the-art comparison on CAMO-FS dataset among the baseline model of MTFA [49], Mask RCNN [66], iFS-RCNN [79], and our proposed methods FS-CDIS with instance triplet loss (ITL) and instance memory storage (IMS). Our performance improves over the utilized baselines.
Model Novel AP Method Baseline Instance Segmentation Object Detection 1 2 3 5 Avg. 1 2 3 5 Avg. MTFA [49] COCO-80 ResNet-50 2.48 6.67 5.81 6.40 5.34 1.98 6.47 5.82 6.17 5.11 M-RCNN [66] 4.08 6.79 6.90 8.29 6.52 2.82 5.09 5.46 6.18 4.89 iFS-RCNN [79] 4.17 6.26 5.73 6.38 5.64 3.92 6.06 5.47 6.60 5.51 MTFA [49] COCO-80 ResNet-101 3.66 6.21 6.16 5.95 5.50 2.93 5.90 5.84 5.84 5.13 M-RCNN [66] 4.39 7.69 7.94 10.09 7.53 3.03 5.80 6.20 7.79 5.71 iFS-RCNN [79] 4.27 6.55 6.07 7.80 6.17 3.79 6.28 6.01 8.08 6.04 Our performance FS-CDIS-ITL 4.46 5.57 6.41 8.48 6.23 4.04 7.28 7.49 9.76 7.14 FS-CDIS-IMS ResNet-101 MTFA 5.46 6.95 7.36 9.61 7.35 4.50 6.95 7.55 10.36 7.34 FS-CDIS-ITL 5.73 7.97 8.52 9.92 8.04 5.08 7.56 7.85 9.67 7.34 FS-CDIS-IMS ResNet-101 M-RCNN 5.52 7.84 8.65 9.82 7.96 4.92 7.39 7.96 9.52 7.45 FS-CDIS-ITL 5.35 6.01 7.80 9.35 7.13 4.71 5.66 7.10 10.36 6.96 FS-CDIS-IMS ResNet-101 iFS-RCNN 2.99 6.83 6.14 9.03 6.25 2.74 6.39 5.94 8.44 5.88 M-RCNN is Mask R-CNN [66] with sigmoid classifier.